# Decoupling Strategy and Execution in Task-Focused Dialogue via Goal-Oriented Preference Optimization

### v1

@@ 1 Introduction
Figure 1: TSE and Reward Across Datasets: GOPO-Qwen3-14B vs. Qwen-235B, DeepSeek-R1, and GPT-5.2.
We identify three interrelated challenges underlying this issue. First, optimization objectives are misaligned with business goals. Mainstream approaches such as supervised fine-tuning and RLHF prioritize technical criteria like preference alignment and semantic relevance, while overlooking business outcomes, resulting in fluent yet ineffective responses that fail to drive conversion or resolve complex issues. Second, existing single-agent architectures lack strategic flexibility in complex multi-turn dialogues. By conflating strategy selection with response generation, they struggle to adapt to dynamic user intentions and emotional shifts, leading to repetitive or fragmented interactions and degraded user experience. Finally, standard operating procedure (SOP) compliance remains unreliable. Prompt-based soft constraints cannot consistently enforce enterprise rules such as mandatory disclaimers or commitment prohibitions, causing strategic intent to be diluted during execution and leading to unstable service quality and compliance risks.
To address these challenges, we propose GOPO, a hierarchical reinforcement learning (RL) framework for task-oriented dialogue. GOPO adopts a dual-agent architecture that decouples strategy planning from response execution, comprising an Expert Agent for high-level decision-making and a Customer Service Agent for compliant response generation. The Expert Agent analyzes dialogue states and dynamically assembles strategy chains from a skill pool, which are translated into structured hard constraints and directly enforced by the Customer Service Agent, ensuring strict alignment between strategic intent and generated responses. Moreover, GOPO employs a joint reward model tied to business end-state metrics, aligning long-horizon optimization with real business value.
The primary contributions of this work are summarized as follows:
@@
The remainder of this paper is organized as follows. Section 2 reviews the related work. Section 3 presents the methodology of the GOPO framework in detail. Section 4 presents the experimental setup, results, and an in-depth analysis of the experimental findings. Finally, Section 5 concludes the paper and outlines directions for future research.
@@ 2.1 Agent Architectures in Task-Oriented Dialogue Systems
Task-oriented dialogue systems in e-commerce are typical carriers of business value orientation(Sanjaya et al., 2023; Konda et al., 2021; Gamboa-Cruzado et al., 2023). Their core goal is to achieve a closed loop of demand matching, trust building and order conversion via multi-turn interactions, making this a major research focus recently(Kanojiya et al., 2021). Existing studies on e-commerce dialogue systems fall broadly into two categories: traditional rule-based or modular pipeline methods, and LLM-based end-to-end single-agent approaches(Puertas et al., 2024). Traditional systems adopt a modular pipeline architecture, sequentially executing natural language understanding (NLU), dialogue management (DM) and natural language generation (NLG). With the rise of LLMs, however, end-to-end single-agent architectures have become dominant. Represented by ReAct(Wei et al., 2025) and its variants, these architectures use LLMs as central controllers and implement a “think–act” loop for complex reasoning and tool-use tasks.
Nevertheless, these systems face limitations in complex, long-horizon business processes. A single agent must both select SOP-driven strategies and generate compliant responses, tightly coupling decision logic with execution constraints. Under uncertainties such as ambiguous user intent or emotional fluctuations, this leads to suboptimal strategies, compliance failures, or dialogue stagnation.
To overcome single-agent limitations, the community has explored multi-agent approaches. CAMEL(Li et al., 2023) uses a role-playing framework where two agents debate solutions, while other works design agent teams with explicit role divisions, e.g., a router agent dispatching requests. Although these methods partially decompose tasks, they rely on natural language soft constraints for inter-agent communication, which is unreliable in strict SOP scenarios, as LLM stochasticity can dilute strategic intent. In contrast, GOPO employs a hierarchical dual-agent architecture with structured hard-constraint transmission, ensuring Expert Agent’s strategies are executed by the Customer Service Agent.
@@ 2.2 Optimization Methods for Dialogue Policies
RL has long been a core approach for optimizing dialogue policies (Peiyuan et al., 2024). Classical methods such as Proximal Policy Optimization (PPO) maximize cumulative rewards, but in multi-turn dialogues rewards are sparse and delayed, with transaction success only observable at the end of a conversation. This leads to severe long-horizon credit assignment issues, making optimization inefficient and prone to local optima, such as over-emphasizing fluency while neglecting SOP compliance.
To mitigate reward design challenges, preference-based optimization methods, particularly Direct Preference Optimization (DPO) and its variants, have gained attention. DPO leverages pairwise human preferences to avoid explicit reward modeling and online RL, and subsequent frameworks (Wu et al., 2025) extend this paradigm to real-time dialogue settings using large-scale offline data. Building on this line of work, Single-stream Policy Optimization (SPO) (Xu and Ding, 2025) and Stepwise Direct Preference Optimization (SDPO) (Kim et al., 2025) introduce finer-grained supervision. However, these approaches rely on static preferences and primarily optimize technical alignment metrics, without explicitly incorporating outcome-driven business signals such as user satisfaction or purchase conversion. Recent methods such as (Zhou et al., 2025) further explore trajectory-level optimization but remain focused on generic preference consistency rather than task-specific business objectives. In contrast, GOPO adopts a joint reward model directly tied to business metrics and, through a hierarchical RL design, provides timely signals that effectively address long-horizon credit assignment.
@@ 2.3 SOP Compliance and Constraint Enforcement
In heavily regulated industries such as finance, healthcare and e-commerce, strict adherence to standard operating procedures is critical for dialogue systems. Existing methods enforce constraints mainly via prompt engineering or rule-based post-processing filtering. However, prompt-based ”soft constraints” are fragile against LLMs’ generative creativity and cannot ensure full compliance [16]. Post-processing filters are passive remedial measures that fail to fundamentally prevent non-compliant outputs and may impair response fluency and naturalness.
Some studies attempt to introduce stronger constraints during decoding. For example, constrained beam search(Chen and Authors, 2025) or vocabulary masking during decoding can enforce the generation of specific keywords or prevent prohibited terms. While these methods improve compliance to some extent, they generally struggle with complex logical constraints and may conflict with the model’s internal knowledge, leading to degraded generation quality.
Our GOPO framework addresses this issue through a novel hard-constraint transmission mechanism. The Expert Agent not only decides the SOP strategy but also compiles it into a set of structured hard constraints, which are directly integrated into the Customer Service Agent’s generation process. Additionally, we design an SOP matching reward module that provides real-time validation and rewards for compliance at the execution layer, creating a complete closed loop from strategy decision-making to response generation and ensuring strict adherence to SOPs.
@@ 3.1 Framework Overview
Figure 2 presents the hierarchical dual-agent architecture of GOPO. The architecture consists of an Expert Agent for strategy planning and a Customer Service Agent for constrained response generation. Under a HRL paradigm, the Expert Agent composes skills from a predefined pool, decoupling policy planning from content generation.
@@
This hierarchical division decomposes dialogue tasks into two coupled subproblems, allowing independent optimization of strategy and response generation.
@@ 3.2 Problem Formalization
  • Action ($a_{t}^{A}\in A_{A}$): The action space of $A$. $A^{A}$ is continuous, representing the generated response text sequence. $a_{t}^{A}=(w_{1},w_{2},\dots,w_{m})$, where $w_{i}$ is a token from the vocabulary.
@@
Based on the MDP formulation, GOPO aims to select skills to define turn-level strategies and generate constrained responses that satisfy immediate dialogue needs while maximizing long-term business objectives under strict compliance.
Table 1: Comparison of GOPO (Qwen-7B-Chat) with Large-Scale Baselines Across Multiple Datasets (TSE, Reward, BLEU).
@@
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
@@
Table 2: Comparison of GOPO with Large-Scale Baselines Across Multiple Datasets (TSE, Reward, BLEU).
@@
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
@@
$\max_{\pi_{E},\pi_{A}}\;\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R_{\text{total}}(s_{t},a_{t}^{E},a_{t}^{A})\right].$ (1)
3.3 Reward and Loss Design
The Expert Agent reward $R_{t}^{E}$ evaluates skill selection quality via Expert Skill Normalized Discounted Cumulative Gain (ESNDCG), which measures the relevance between the predicted skill sequence $P=[p_{1},\dots,p_{m}]$ and the Teacher-generated reference sequence $T=[t_{1},\dots,t_{n}]$. For any predicted skill $s$, its relevance is defined as:
Based on this relevance definition, we compute the Discounted Cumulative Gain (DCG) for the predicted skill sequence $P=[p_{1},p_{2},\dots,p_{m}]$ as:
$\text{DCG}(P)=\sum_{i=1}^{m}\frac{2^{\text{Rel}(p_{i})}-1}{\log_{2}(i+1)}.$ (3)
which represents the theoretical maximum score. The normalized metric, Expert Skill Normalized Discounted Cumulative Gain (ESNDCG)(Zhang et al., 2022), is then computed as
$\text{ESNDCG}(P,T)=\frac{\text{DCG}(P)}{\text{IDCG}(T)}.$ (5)
Finally, the reward for $E$ is defined as:
$R_{t}^{E}=\text{ESNDCG}(P,T).$ (6)
$R_{t}^{A}=\alpha_{1}S_{1}+\alpha_{2}S_{2}+\alpha_{3}S_{3}+\alpha_{4}S_{4},$ (8)
where $\alpha_{i}$ are normalized weights controlling the contribution of each evaluation dimension.
$\mathcal{L}_{t,\text{s}}(\theta_{A})$ $=d(a_{t}^{A},a_{t}^{E}),$ (9c)
@@ 3.3 Reward and Loss Design
Table 4: Comparison of GOPO Variants on Multiple Datasets (TSE, Reward, BLEU).
@@
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
@@
$R_{t}=w_{t}^{E}\cdot R_{t}^{E}+w_{t}^{A}\cdot R_{t}^{A}.$ (10)
Here, the Expert Agent weight $w_{t}^{E}$ increases over turns, shifting emphasis from early response quality to later strategic skill selection for business goals.
@@
4 Experiments
4.1 Experimental Setup
Model Mgshop TmallBrand-A
TSE Reward BLEU TSE Reward BLEU
Full GOPO 0.9475 7.63 0.279 0.8511 7.46 0.186
GOPO w/o Expert Agent 0.9371 7.16 0.217 0.8023 6.47 0.227
Full Baseline 0.7353 6.92 0.251 0.6371 6.43 0.213
Baseline w/o Expert Agent 0.4327 6.08 0.240 0.1818 5.19 0.182
Table 6: Performance of Customer Service Agents Evaluated with Different Expert Agents on Mgshop Dataset.
Expert Agent/ Customer Service Agent GOPO-Qwen3-14B Qwen-235B DeepSeek-R1 Gemini-2.5
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
GOPO-Qwen3-14B 0.9475 7.63 0.279 0.8714 7.40 0.193 0.7854 6.73 0.234 0.7384 5.30 0.251
Qwen-235B 0.9352 7.45 0.214 0.9239 7.37 0.186 - - - - - -
DeepSeek-R1 0.9453 7.55 0.242 - - - 0.9426 7.51 0.145 - - -
Gemini-2.5 0.9268 7.52 0.220 - - - - - - 0.9265 7.12 0.130
Table 7: Performance of Customer Service Agents Evaluated with Different Expert Agents on TmallBrand-A Dataset.
Expert Agent/ Customer Service Agent GOPO-Qwen3-14B Qwen-235B DeepSeek-R1 Gemini-2.5
TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU TSE Reward BLEU
GOPO-Qwen3-14B 0.8511 7.46 0.186 0.8291 6.27 0.203 0.8395 6.84 0.146 0.8163 6.55 0.228
Qwen-235B 0.8428 7.37 0.201 0.8371 6.54 0.235 - - - - - -
DeepSeek-R1 0.8503 7.03 0.245 - - - 0.8443 7.50 0.218 - - -
Gemini-2.5 0.8025 6.83 0.131 - - - - - - 0.8152 6.87 0.196
4.1.1 Datasets
  • Mgshop: An internal test set designed to evaluate robustness under highly dynamic user emotions and rapidly changing intents.
@@ 4.1.2 Evaluation Metrics
  • Task-focused Sequential Engagement Success (TSE): A sequence-level metric that measures task completion efficiency in Task-focused dialogues, covering requirement matching, purchase-stimulating information delivery, and proactive guidance. The mathematical formula for TSE is given as follows:
@@
    Here, $\omega_{1},\omega_{2},\omega_{3}$ are task weights ($\sum\omega_{i}=1$), $I_{i}$ indicate task completion, $\gamma$ is a turn-based decay factor, and $N_{i}$ denote the turns at which each task is completed. TSE directly measures the effectiveness of business decision-making by assessing whether key tasks are completed in minimal dialogue turns.
  • Reward (LM Score): Responses are scored by a large model (e.g., GPT-4) and averaged over dialogue turns.
  • BLEU(Papineni et al., 2002): Evaluates n-gram overlap with references but lacks semantic sensitivity.
  • G-Eval(Liu et al., 2023): An LLM-based metric assessing coherence, completeness, and emotional appropriateness via structured prompts.

### v2

@@ 1 Introduction
Figure 1: TSE, GRE, and G-Eval Across Datasets: GOPO-Qwen3-14B vs. Qwen-235B, DeepSeek-R1, and GPT-5.2.
We identify three interrelated challenges underlying this issue. First, optimization objectives are misaligned with business goals. Mainstream approaches such as supervised fine-tuning and RLHF prioritize technical criteria like preference alignment and semantic relevance, while overlooking business outcomes, resulting in fluent yet ineffective responses that fail to drive conversion or resolve complex issues (Ouyang et al., 2022). Second, existing single-agent architectures lack strategic flexibility in complex multi-turn dialogues (Wang et al., 2025). By conflating strategy selection with response generation, they struggle to adapt to dynamic user intentions and emotional shifts, leading to repetitive or fragmented interactions and degraded user experience. Finally, standard operating procedure (SOP) compliance remains unreliable. Prompt-based soft constraints cannot consistently enforce enterprise rules such as mandatory disclaimers or commitment prohibitions, causing strategic intent to be diluted during execution and leading to unstable service quality and compliance risks.
To address these challenges, we propose GOPO, a hierarchical reinforcement learning (RL) framework for task-oriented dialogue. GOPO adopts a dual-agent architecture that decouples strategy planning from response execution, comprising an Expert Agent for high-level decision-making and a Customer Service Agent for compliant response generation. The Expert Agent analyzes dialogue states and dynamically composes strategy chains from a skill pool, which are translated into structured hard constraints and enforced by the Customer Service Agent to align strategic intent with generated responses across multi-turn interactions. Moreover, GOPO employs a joint reward model grounded in business end-state metrics, aligning long-horizon optimization with real business value. The primary contributions of this work are summarized as follows:
@@
The remainder of this paper is organized as follows. Section 2 reviews the related works. Section 3 presents the methodology of the GOPO framework in detail. Section 4 presents the experimental setup, results, and an in-depth analysis of the experimental findings. Finally, Section 5 concludes the paper and outlines directions for future research.
@@ 2.1 Agent Architectures in Task-Oriented Dialogue Systems
Task-oriented dialogue systems in e-commerce serve as carriers of business value (Sanjaya et al., 2023; Konda et al., 2021; Gamboa-Cruzado et al., 2023). Their core objective is to establish a closed loop of demand matching, trust building, and order conversion through multi-turn interactions, making this a central research focus (Kanojiya et al., 2021). Existing approaches broadly fall into two categories: traditional rule-based or modular pipeline systems, and LLM-based end-to-end single-agent methods (Puertas et al., 2024). The former rely on sequential NLU–DM–NLG pipelines, while the latter exemplified by ReAct (Wei et al., 2025) and its variants use LLMs as central controllers to implement a “think–act” loop for complex reasoning and tool use (Yao et al., 2022).
Nevertheless, complex long-horizon business processes expose key limitations. A single agent must jointly handle SOP-driven strategy selection and compliant response generation, coupling decision logic with execution. Under uncertainty, this often results in suboptimal strategies, compliance failures, or stalled dialogues. To overcome single-agent limitations, the community has explored multi-agent approaches. CAMEL (Li et al., 2023) used a role-playing framework where two agents debate solutions, while other works designed agent teams with explicit role divisions, e.g., a router agent dispatching requests. Some work achieves zero-shot multi-agent coordination by training agents with a single partner across environments to learn general cooperative norms (Jha et al., 2025). Although these methods partially decompose tasks, they relied on natural language soft constraints for inter-agent communication, which was unreliable in strict SOP scenarios, as LLM stochasticity can dilute strategic intent. In contrast, GOPO employs a hierarchical dual-agent architecture with structured hard-constraint transmission, ensuring Expert Agent’s strategies are executed by the Customer Service Agent.
@@ 2.2 Optimization Methods for Dialogue Policies
RL has long been a core approach for optimizing dialogue policies (Feng et al., 2024; Wu et al., 2025). Classical methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) maximized cumulative rewards, but in multi-turn dialogues rewards were sparse and delayed, with transaction success only observabled at the end of a conversation (Liu et al., 2025). This led to severe long-horizon credit assignment issues, making optimization inefficient and prone to local optima, such as over-emphasizing fluency while neglecting SOP compliance.
To mitigate reward design challenges, preference-based optimization methods—particularly Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants have gained attention. DPO leveraged pairwise human preferences to bypass explicit reward modeling and online RL, and subsequent work extended this paradigm to real-time dialogue using large-scale offline data. Building on this line, Single-stream Policy Optimization (SPO) (Xu and Ding, 2025) and Stepwise Direct Preference Optimization (SDPO) (Kim et al., 2025a) introduced finer-grained supervision. However, these approaches relied on static preferences and primarily optimize technical alignment metrics, without explicitly incorporating outcome-driven business signals such as user satisfaction or purchase conversion.
Recent methods further explored long-horizon optimization by introducing trajectory-level or step-level signals. Memento (Zhou et al., 2025a) focused on trajectory-level preference consistency, while COLLABLLM (Li et al., 2025) employed multiturn-aware rewards to encourage proactive intent discovery. SWEET-RL (Zhou et al., 2025b) addressed multi-turn credit assignment by training a critic with access to training-time information to provide step-level rewards. In contrast, GOPO adopts a joint reward model directly tied to business metrics and, through a hierarchical RL design, delivers timely supervision that effectively addresses long-horizon credit assignment.
@@ 2.3 SOP Compliance and Constraint Enforcement
In heavily regulated industries such as finance, healthcare, and e-commerce, strict adherence to standard operating procedures (SOPs) is critical for dialogue systems (Gao et al., 2018). Existing approaches mainly enforced constraints through prompt engineering or rule-based post-processing. However, prompt-based “soft constraints” are fragile to the generative flexibility of LLMs and cannot reliably ensure compliance. Post-processing filters act as passive remedies that fail to fundamentally prevent non-compliant outputs and may degrade response fluency and naturalness.
Some studies introduced stronger constraints during decoding. For instance, constrained beam search (Chen and others, 2025) and vocabulary masking (Kim et al., 2025b) can enforce required keywords or block prohibited terms. While these techniques improved compliance to a limited extent, they struggled with complex logical constraints and may conflict with the model’s internal representations, resulting in degraded generation quality.
GOPO addresses these limitations through a hard-constraint transmission mechanism. The Expert Agent not only determines SOP-level strategies but also compiles them into structured hard constraints that are integrated into the Customer Service Agent’s generation process. In addition, an SOP-matching reward module provides real-time compliance feedback at the execution level, forming a closed loop from strategy planning to response generation and enabling consistent adherence to SOPs.
@@ 3.1 Framework Overview
Figure 2 presents the hierarchical dual-agent architecture of GOPO. The architecture consists of an Expert Agent for strategy planning and a Customer Service Agent for constrained response generation. Under a hierarchical RL paradigm, the Expert Agent composes skills from a predefined pool, decoupling policy planning from content generation.
@@
Despite introducing hierarchical dependence, this design reduces exploration complexity and gradient variance, enabling stable long-horizon optimization.
@@ 3.2 Problem Formalization
  • Action ($a_{t}^{A}\in A_{A}$): The action space of $A$. $A^{A}$ is continuous, representing the generated response text sequence. $a_{t}^{A}=(w_{1},w_{2},\dots,w_{m})$, where $w_{i}$ is a token from the vocabulary $V$.
@@
Based on the MDP formulation, GOPO selects skills to define strategies for each dialogue turn and generates responses that satisfy immediate dialogue requirements while ensuring strict compliance and maximizing long-term business objectives.
$\max_{\pi_{E},\pi_{A}}\;\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R_{t}(s_{t}^{E},s_{t}^{A},a_{t}^{E},a_{t}^{A})\right].$ (1)
3.3 Reward and Loss Design
The Expert Agent reward $R_{t}^{E}$ evaluates the quality of skill selection by measuring the relevance between a predicted skill sequence $P=[p_{1},\dots,p_{m}]$ and a Teacher-generated reference sequence $T=[t_{1},\dots,t_{n}]$. The teacher reference sequence is generated offline and serves as a stable preference anchor during training, avoiding the need for real-time oracle supervision at deployment. To quantify sequence-level relevance, we build upon Discounted Cumulative Gain (DCG) (Qiu et al., 2022), a standard ranking-based metric that accounts for both relevance and position. For any predicted skill $s$, its relevance is defined as:
Based on this relevance definition, the DCG of the predicted skill sequence $P=[p_{1},p_{2},\dots,p_{m}]$ is computed as:
$\text{DCG}(P)=\sum_{i=1}^{m}\frac{2^{\text{Rel}(p_{i})}-1}{\log_{2}(i+1)}.$ (3)
which represents the theoretical maximum score. We then define Expert Skill Normalized Discounted Cumulative Gain (ESNDCG) as a normalized variant of DCG, which is then computed as
$\text{ESNDCG}(P,T)=\frac{\text{DCG}(P)}{\text{IDCG}(T)}.$ (5)
Finally, the reward for $E$ is defined as:
$R_{t}^{E}=\text{ESNDCG}(P,T).$ (6)
Although ESNDCG relies on teacher-generated references for reward shaping, it provides a trajectory-level preference signal rather than a supervised target, with optimization carried out via policy gradients. In practice, both the predicted and teacher skill sequence lengths are capped at 5, ensuring numerical stability while emphasizing short-horizon, high-impact skill planning.
$R_{t}^{A}=\alpha_{1}S_{1}+\alpha_{2}S_{2}+\alpha_{3}S_{3}+\alpha_{4}S_{4},$ (8)
where $\alpha_{i}$ are normalized weights controlling the contribution of each evaluation dimension, and $S_{i}$ represent the individual scores for each evaluation dimension.
$\mathcal{L}_{t,\text{s}}(\theta_{A})$ $=d(a_{t}^{A},a_{t}^{E}),$ (9c)
$R_{t}=w_{t}^{E}\cdot R_{t}^{E}+w_{t}^{A}\cdot R_{t}^{A}.$ (10)
Here, the Expert Agent weight $w_{t}^{E}$ increases over turns, shifting emphasis from early response quality to later strategic skill selection for business goals. We adopt a temporally adaptive scheme where $w_{t}^{E}$ increases over turns while $w_{t}^{A}$ remains strictly positive, continuously enforcing response-level quality and politeness constraints.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
Table 1: Comparison of GOPO (Qwen-7B-Chat) with Large-Scale Baselines Across Multiple Datasets (TSE, GRE, BLEU).
@@ 4.1.1 Datasets
TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU
@@
Table 2: Comparison of GOPO with Large-Scale Baselines Across Multiple Datasets (TSE, GRE, BLEU).
@@
TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU
@@
Table 4: Comparison of GOPO Variants on Multiple Datasets (TSE, GRE, BLEU).
@@
TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU TSE GRE BLEU
@@
  • Mgshop (Jannach et al., 2021): An internal test set designed to evaluate robustness under highly dynamic user emotions and rapidly changing intents.
@@ 4.1.2 Evaluation Metrics
  • Task-focused Sequential Engagement Success (TSE): A sequence-level metric measuring task completion efficiency in task-focused dialogues, covering requirement matching, information delivery, and proactive guidance. The TSE score is defined as:
@@
    Here, $\omega_{i}$ are task weights, $I_{i}$ indicate task completion, $\gamma$ is a turn-based decay factor, and $N_{i}$ denote task completion turns. TSE measures whether key tasks are completed efficiently.
  • Goal-oriented Response Effectiveness (GRE): A response-level metric evaluating how individual replies advance business objectives, computed by averaging LLM-based turn-level scores on politeness, product appropriateness, and action guidance.
  • BLEU (Papineni et al., 2002): Measures n-gram overlap with reference responses, capturing surface-level lexical similarity.
  • G-Eval (Liu et al., 2023): An LLM-based metric assessing dialogue fluency and user experience via coherence, completeness, and emotional appropriateness.
Overall, these metrics capture complementary aspects of performance across multiple levels, ranging from sequence-level goal achievement to response-level goal advancement, user experience, and surface-level similarity.